1) A diffraction grating is w-3 cm wide and has N-1.2 x 104 slits. A monochromatic light source of wavelength λ-680 nm shines through the grating onto a screen. At what angle does the first-order line appear on the screen? 8-1 5.22 deg 8-1 5.78 deg θ-105.32 deg θ-59.54 deg 0 143.24 deg Submit 2) Another source of light is present and is just barely resolved from the 680 nm source in the second order. What is the difference in wavelength between this source and the original source? Αλ-0.014 nm Αλ-0.028 nm 0.042 nm 0.057 nm -1.039 nm Submit 3) If N is doubled, what happens to the minimum difference in wavelength that can be resolved for a given order? It decreases Οκ It remains the same Οκ It increases Submit3) If N is doubled, what happens to the minimum difference in wavelength that can be resolved for a given order? It decreases It remains the same It increases 4) A laser is directed at a double-slit arrangement. If we increase the wavelength of the light, which of the following will occur to the pattern on a distant screen? The spacing between the interference fringes will decrease, but the width of the overall diffraction envelope will The spacing between the interference fringes will increase, but the width of the overall diffraction envelope will Both the spacing between the interference fringes and the width of the overall diffraction envelopewill increase increase decrease Submit 5) Considering larger objects, a skyscraper has two tall antennae separated by 3 meters. Each antenna has one blue and one red light. You can resolve the two blue lights but not the red ones Assume that your pupils have a diameter = 2 mm at night, and that blue lights wavelength is 475 nm, while red lights wavelength is 650 nm Approximately how far are you from the skyscraper? (Also assume that you have perfect vision limited only by diffraction.) 5 km 10 km 15 km Submit
Answer #1

1)We know that

d sin(theta) = m lambda

theta = sin^-1 (m lambda/d)

d = W/N = 3/(1.2 x 10^4)

theta = sin^-1 (1 x 680 x 10^-12/0.0000025) = 15.78 deg

Hence, theta = 15.78 deg

-----------------

2)delta-lambda(min) = lambda/(N m)

delta-lambda(min) = 680/(1.2 x 10^4 x 2) = 0.028

Hence, delta-lambda = 0.028

-------------------

decreased

(looking at the above relation used in 2, it decreases)

------------------